the sum of the digits of a two-digit number is 14. if the digits are reversed, the new number is 18 less than the original number. find original number
Answer by palanisamy(496)

Let the number be xy
Given, x+y = 14 ...(1)
Also,if the digits are reversed, the new number is 18 less than the original number.
10y+x = 10x+y-18
10y+x-10x-y = -18
9y-9x = -18
y-x = -2 ...(2)
(1)+(2) => 2y = 12
y = 6
(1)=> x+6 = 14
x = 14-6=8
So the required number is 86